# Integrating Weather Station Data and Radar for Precipitation Nowcasting: SmaAt-fUsion and SmaAt-Krige-GNet

### v1

@@ 3.4 SmaAt-Krige-GNet model
In this work, we explicitly integrate spatial dependencies between weather stations into the SmaAt-UNet model. To achieve this, we first generate grid-based maps of station variables using Kriging [28, 29]. Inspired by [23], which introduced SmaAt-GNet as a generator within an adversarial learning framework for precipitation nowcasting, we propose SmaAt-Krige-GNet. This model combines the grid-based maps derived from station data with precipitation maps. The key distinction between SmaAt-GNet and the standard SmaAt-UNet model lies in the addition of an extra encoder stream. This new stream mirrors the encoder layers of SmaAt-UNet and concatenates its representations at each level with those of the original SmaAt-UNet encoder. While the original SmaAt-GNet in [23] was designed to integrate binary precipitation masks into the SmaAt-UNet model, our proposed SmaAt-Krige-GNet instead utilizes Kriging-based spatial maps. By structuring the model into two independent streams rather than stacking all inputs channel-wise in a single-stream architecture, SmaAt-Krige-GNet can independently learn hidden features from both weather and precipitation data. Since all models process 12 input images, we modify the secondary input of SmaAt-GNet to also accept 12 images. Given that there are 8 weather variables, the resulting input forms a 4D tensor of size $64\times 64\times 8\times 12$, which is then flattened into a $64\times 64\times 96$ layer, as illustrated in Fig. 3(b). To explicitly incorporate spatial data, Kriging is applied to interpolate weather station measurements into 2D geospatial maps, ensuring that each weather variable at each timestep is represented in the same spatial resolution as the precipitation maps.
@@ 4.1 Model setup
The models were trained on the dataset for a maximum of 200 epochs, with a batch size of 16. These hyperparameters were chosen based on preliminary experimentation to balance model performance and computational efficiency. The implementation of our models is available at GitHub  [^2] https://github.com/JieJieNiu/SmaAt-fUsion-and-SmaAt-Krige-GNet.
To generate the Kriging maps for SmaAt-Krige-GNet, we use Ordinary Kriging from PyKrige [30]. A grid is defined with the same resolution as the precipitation maps, specifically $64\times 64$. Subsequently, a variogram is fitted using the spherical model, which has been shown to be suitable for meteorological applications [31, 32]. To ensure accurate spatial representation, we specify the coordinate type as geographic, allowing the model to directly use the latitude and longitude of each weather station. However, Kriging with PyKrige can be numerically unstable when using its default inverse matrix algorithm. To mitigate this, we set the ‘pseudo_inv’ parameter of Ordinary Kriging to True, which stabilizes computations at the cost of increased processing time [30]. Once the linear system is solved, a set of weights $\lambda_{i}$ is obtained, which are then applied to interpolate the Kriging value at each coordinate within the grid. This process generates spatially continuous weather variable maps that align with the resolution of the precipitation data.
Figure 4: Example Kriging map input of each variable
@@ 4.2 Evaluation
The loss function utilized in this study is the Mean Squared Error (MSE), which quantifies the difference between the predicted and ground truth images. In addition to MSE, and in line with previous works [23, 10, 33, 34], we employ a set of widely used binary metrics for precipitation nowcasting, including the F1-Score, Critical Success Index (CSI), Heidke Skill Score (HSS), and Matthews Correlation Coefficient (MCC), to assess the performance of the proposed models. Since precipitation maps contain continuous values, they must be thresholded to convert them into binary masks before applying these metrics. In this study, we evaluate precipitation at three different thresholds: 0.5 mm/h, 10 mm/h, and 20 mm/h. Applying a threshold to the predicted and target images produces a binary mask, where each pixel is assigned 1 if its value exceeds the threshold (rain) and 0 otherwise (no rain). From this, one can calculate the true positives (TP) (prediction=1, target=1), false positives (FP) (prediction=1, target=0), true negatives (TN) (prediction=0, target=0) and false negatives (FN) (prediction=0, target=1). Using thresholds, the models’ performance can be assessed across varying precipitation intensities. The 0.5 mm/h threshold is commonly used as it effectively captures light rain [35]. Additionally, 10 mm/h and 20 mm/h thresholds are tested to evaluate the models’ ability to predict moderate to extreme precipitation events. The binary metrics CSI, HSS and MCC are computed as follows:
@@
The CSI evaluates the overlap between predicted and observed precipitation events [36]. It accounts for false positives and missed events, providing a robust measure for precipitation predictions [37]. The HSS compares the model predictions to a random baseline [37]. HSS highlights the skill of the model beyond chance-level prediction [38]. Unlike accuracy, which can be misleading in imbalanced datasets, the MCC score provides a more reliable measure by considering both classes (positive and negative). The MCC ranges from -1 (perfect inverse correlation) to +1 (perfect correlation), with 0 indicating no better than random predictions. This metric is particularly useful when the data is highly imbalanced, as it accounts for both the correct and incorrect predictions across all categories [39]. In the case of precipitation nowcasting, the present imbalance is the high amount of negative precipitation events compared to positive precipitation events.
@@ 5 Results and Discussion
Figure 5: Example precipitation input of the dataset on 2019-12-14.
Figure 6: Example outputs of each model, using the input seen in Fig. 5
Figure 7: The same outputs as seen in Fig. 6, after applying a minimum threshold of 0.5mm/hour. The scaling has been changed compared to Fig. 6 for improved visibility. The outputs still represent a 5 minute interval.
After training, models were selected based on the lowest validation loss to ensure optimal generalization. The obtained results are presented in Table 3. These results were computed after denormalizing the data, following the approach used in similar studies [23, 10].
@@
At higher rainfall thresholds (10 mm/h and 20 mm/h), SmaAt-Krige-GNet underperforms relative to both SmaAt-fUsion and the core SmaAt-UNet in terms of binary evaluation metrics. This performance decrease is likely due to the model’s tendency to over-smooth predictions, primarily a consequence of incorporating Kriging-generated maps. As shown in Fig. 4, these maps inherently introduce smoothness. Since weather stations are sparsely distributed over a large area, the interpolated values in the Kriging maps may not fully capture the fine-scale variations in actual weather conditions. A potential solution to this limitation would be a significant increase in the number of weather stations, which could improve the resolution and accuracy of the interpolated Kriging maps.
In Fig. 6, an example output of each model can be observed compared to the ground truth. Figure 7 displays the predictions after applying a 0.5 mm/h threshold, with the southeast region closely matching the ground truth in the SmaAt-Krige-GNet model. Interestingly, only the SmaAt-fUsion model captures the slight precipitation in the north, indicating the presence of information missing in SmaAt-UNet and potentially overlooked by SmaAt-Krige-GNet due to smoothing.
@@ 6 Conclusion
This study introduced two novel deep learning models, SmaAt-fUsion and SmaAt-Krige-GNet, designed to improve precipitation nowcasting by integrating meteorological variables from weather stations alongside radar data. The experimental results demonstrate that both models have distinct advantages: SmaAt-Krige-GNet excels in low-intensity rainfall predictions by incorporating weather station data through Kriging-based spatial interpolation, while SmaAt-fUsion outperforms in high-intensity events by directly integrating station data into the network’s bottleneck. These findings highlight the potential of leveraging both discrete weather station data and spatially interpolated variables to improve deep learning-based precipitation forecasting.

### v2

@@ 3.4 SmaAt-Krige-GNet model
In this work, we explicitly integrate spatial dependencies between weather stations into the SmaAt-UNet model. To achieve this, we first generate grid-based maps of station variables using Kriging Isaaks1989; Matheron1963. Inspired by reulen2024ga, which introduced SmaAt-GNet as a generator within an adversarial learning framework for precipitation nowcasting, we propose SmaAt-Krige-GNet. This model combines the grid-based maps derived from station data with precipitation maps. The key distinction between SmaAt-GNet and the standard SmaAt-UNet model lies in the addition of an extra encoder stream. This new stream mirrors the encoder layers of SmaAt-UNet and concatenates its representations at each level with those of the original SmaAt-UNet encoder. While the original SmaAt-GNet in reulen2024ga was designed to integrate binary precipitation masks into the SmaAt-UNet model, our proposed SmaAt-Krige-GNet instead utilizes Kriging-based spatial maps. By structuring the model into two independent streams rather than stacking all inputs channel-wise in a single-stream architecture, SmaAt-Krige-GNet can independently learn hidden features from both weather and precipitation data. Since all models process 12 input images, we modify the secondary input of SmaAt-GNet to also accept 12 images. Given that there are 8 weather variables, the resulting input forms a 4D tensor of size $64\times 64\times 8\times 12$, which is then flattened into a $64\times 64\times 96$ layer, as illustrated in Fig. 3(b). To explicitly incorporate spatial data, Kriging is applied to interpolate weather station measurements into 2D geospatial maps, ensuring that each weather variable at each timestep is represented in the same spatial resolution as the precipitation maps.
@@
Table 4: Ablation study for the eight meteorological variables in SmaAt-Krige-GNet and SmaAt-fUsion at threshold $\geq 0.5\text{ mm/h}$. Removing each variable reveals its relative contribution to precipitation nowcasting performance.
[…]
SmaAt-Krige-GNet None (all variables) 0.7808 0.6404
Humidity 0.7638 0.0170 2.18% 0.6200 0.0204 3.19%
Atmospheric pressure 0.7639 0.0169 2.16% 0.6222 0.0182 2.84%
Wind direction 0.7646 0.0162 2.07% 0.6222 0.0182 2.84%
Avg. wind speed 0.7650 0.0158 2.02% 0.6226 0.0178 2.78%
Max. wind speed 0.7727 0.0081 1.04% 0.6296 0.0108 1.69%
Std. wind speed 0.7801 0.0007 0.09% 0.6403 0.0001 0.02%
Air temperature 0.7808 0.0000 0.00% 0.6404 0.0000 0.00%
Std. wind direction 0.7808 0.0000 0.00% 0.6404 0.0000 0.00%
SmaAt-fUsion None (all variables) 0.7680 0.6231
Wind direction 0.7535 0.0145 1.89% 0.6088 0.0143 2.29%
Atmospheric pressure 0.7539 0.0141 1.84% 0.6080 0.0151 2.42%
Humidity 0.7552 0.0128 1.67% 0.6074 0.0157 2.52%
Avg. wind speed 0.7572 0.0108 1.41% 0.6062 0.0169 2.71%
Max. wind speed 0.7608 0.0072 0.94% 0.6159 0.0072 1.16%
Air temperature 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
Std. wind direction 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
Std. wind speed 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
[…]
@@ 4.1 Model setup
The models were trained on the dataset for a maximum of 200 epochs, with a batch size of 16. These hyperparameters were chosen based on preliminary experimentation to balance model performance and computational efficiency.
To generate the Kriging maps for SmaAt-Krige-GNet, we use Ordinary Kriging from PyKrige pykrige. A grid is defined with the same resolution as the precipitation maps, specifically $64\times 64$. Subsequently, a variogram is fitted using the spherical model, which has been shown to be suitable for meteorological applications Verworn2011; CSEGRecorderVariograms. To ensure accurate spatial representation, we specify the coordinate type as geographic, allowing the model to directly use the latitude and longitude of each weather station. However, Kriging with PyKrige can be numerically unstable when using its default inverse matrix algorithm. To mitigate this, we set the ‘pseudo_inv’ parameter of Ordinary Kriging to True, which stabilizes computations at the cost of increased processing time pykrige. Once the linear system is solved, a set of weights $\lambda_{i}$ is obtained, which are then applied to interpolate the Kriging value at each coordinate within the grid. This process generates spatially continuous weather variable maps that align with the resolution of the precipitation data.
Figure 4: An example of Kriging map input of the eight meteorological variables at a selected timestamp. To improve visibility, Atmospheric pressure uses a percentile-based color scale due to its narrow dynamic range, while the other variables share a common normalization (2-2 to 2).
4.2 Baselines
To comprehensively evaluate the performance of the proposed models, we compared them against four representative baselines both deep learning and optical flow approaches. The Persistence model simply assumes that future precipitation will remain the same as the latest observation. SmaAt-UNet trebing2021smaat represents a compact attention-based UNet widely adopted in radar nowcasting. In addition, two optical flow methods were included: PySTEPS pulkkinen2019pysteps which performs stochastic extrapolation using motion fields, and RainyMotion ayzel2019optical, a CNN-based flow estimation model. Each nowcasting case consists of 12 input radar frames (one hour at 5-minute intervals) and a single prediction at +30 minutes, consistent with prior UNet-based nowcasting studies. Thus, the temporal frequency of evaluation is also 5 minutes, matching the native radar observation cycle, ensuring a fair comparison across architectures.
@@ 4.3 Evaluation
The loss function utilized in this study is the Mean Squared Error (MSE) hodson2021mean, which quantifies the difference between the predicted and ground truth images. In addition to MSE, and in line with previous works reulen2024ga; trebing2021smaat; stanczyk2021deep; van2023improving, we employ a set of widely used binary metrics for precipitation nowcasting, including the F1-Score naidu2023review, Critical Success Index (CSI) jolliffe2003, Heidke Skill Score (HSS) hogan2012, and Matthews Correlation Coefficient (MCC) mcc, to assess the performance of the proposed models. Since precipitation maps contain continuous values, they must be thresholded to convert them into binary masks before applying these metrics.
In this study, we evaluate precipitation at three different thresholds: 0.5 mm/h, 10 mm/h, and 20 mm/h. Applying a threshold to the predicted and target images produces a binary mask, where each pixel is assigned 1 if its value exceeds the threshold (rain) and 0 otherwise (no rain). From this, one can calculate the true positives (TP) (prediction=1, target=1), false positives (FP) (prediction=1, target=0), true negatives (TN) (prediction=0, target=0) and false negatives (FN) (prediction=0, target=1). To ensure statistical robustness, these contingency counts are first accumulated over all 1,557 nowcasting cases from the full 2019 evaluation period, rather than being computed on a per-sample basis. The final CSI, HSS, and MCC values are then obtained by applying the metric formulas to these aggregated pixel-wise totals, resulting in stable estimates that reflect performance across diverse precipitation events. These metrics also guarantees that the reported scores represent the entire annual distribution of rainfall. In the 2019 test set, this distribution is reflected by 263 winter cases (16.9%), 504 spring cases (32.4%), 342 summer cases (22.0%), and 448 autumn cases (28.8%), demonstrating that the evaluation covers all four seasons and a diverse range of meteorological conditions.
Using thresholds, the models’ performance can be assessed across varying precipitation intensities. The 0.5 mm/h threshold is commonly used as it effectively captures light rain tan2018evaluation. Additionally, 10 mm/h and 20 mm/h thresholds are tested to evaluate the models’ ability to predict moderate to extreme precipitation events. The binary metrics CSI, HSS and MCC are computed as follows:
@@
The CSI evaluates the overlap between predicted and observed precipitation events jolliffe2003. It accounts for false positives and missed events, providing a robust measure for precipitation predictions hogan2012. The HSS compares the model predictions to a random baseline hogan2012. HSS highlights the skill of the model beyond chance-level prediction wilks2011. Unlike accuracy, which can be misleading in imbalanced datasets, the MCC score provides a more reliable measure by considering both classes (positive and negative). The MCC ranges from -1 (perfect inverse correlation) to +1 (perfect correlation), with 0 indicating no better than random predictions. This metric is particularly useful when the data is highly imbalanced, as it accounts for both the correct and incorrect predictions across all categories mcc. In the case of precipitation nowcasting, the present imbalance is the high amount of negative precipitation events compared to positive precipitation events.
Figure 5: Examples precipitation input of the dataset on 2019-03-10, 2019-06-15, 2019-09-29, 2019-12-14.
Figure 6: Example outputs of each model, using the input seen in Fig. 5
@@ 5 Results and Discussion
5.1 Model performance
After training, models were selected based on the lowest validation loss to ensure optimal generalization. The obtained results are presented in Table 3. These results were computed after denormalizing the data, following the approach used in similar studies reulen2024ga; trebing2021smaat.
@@ 5.1 Model performance
At higher rainfall thresholds (10 mm/h and 20 mm/h), SmaAt-Krige-GNet underperforms relative to both SmaAt-fUsion and the core SmaAt-UNet in terms of binary evaluation metrics. This performance decrease is likely due to the model’s tendency to over-smooth predictions, primarily a consequence of incorporating Kriging-generated maps. As shown in Fig. 4, these maps inherently introduce smoothness. Although Kriging provides spatial continuity required by convolutional architectures, the limited number of station observations reduces its ability to reconstruct fine-scale atmospheric structures. This smoothing effect is especially noticeable in variables such as atmospheric pressure and wind-related fields, where local gradients are reduced. Consequently, SmaAt-Krige-GNet benefits from these maps in low-intensity rainfall scenarios, where larger-scale patterns dominate but struggle with sharp precipitation peaks, which require higher spatial resolution. Increasing the density of weather stations would directly improve the fidelity of the interpolated fields. A potential solution to this limitation would be a significant increase in the number of weather stations, which could improve the resolution and accuracy of the interpolated Kriging maps.
Fig. 5 and Fig. 6 further clarifies the strengths and remaining limitations of the proposed architectures. Fig. 5 illustrates four representative precipitation maps across different seasons. These inputs contains both stratiform and convective rainfall, providing a challenging test for models relying solely on reflectivity patterns. 2019-03-10, 2019-09-24, and 2019-12-14 correspond to stratiform precipitation events associated with large-scale frontal systems, characterized by broad rainfall distributions. In contrast, 2019-06-15 represents a convective summer case with sharp spatial gradients. Including both regimes enables a comprehensive evaluation of the models’ ability to handle diverse rainfall patterns.
In Fig. 6, most of models correctly capture the general location and shape of the main precipitation areas. However, SmaAt-fUsion and SmaAt-Krige-GNet produce blurred predictions, particularly along convective boundaries. This is most apparent in the summer case (2019-06-15), where sharp convective streaks are substantially smoothed out. By contrast, PySTEPS and RainyMotion do not suffer from such over-smoothing. Their outputs preserve sharper gradients due to the nature of optical-flow extrapolation. Nevertheless, their predictive skill is limited. Both methods show displacement errors, misplacing the precipitation band in several cases, and in some situations their performance is nearly indistinguishable with the Persistence baseline. These results highlight the trade-off between spatial sharpness and predictive accuracy among different nowcasting approaches.
Recent generative nowcasting models such as DGMR ravuri2021skilful, DiffCast yu2024diffcast, and NowcastNet Zhang2023 produce sharper and more realistic predictions, particularly at longer lead times. These models also achieve strong performance without requiring additional meteorological inputs. In contrast, our approach focuses on multi-data integration by incorporating weather-station variables, which strengthens the physical basis of the predictions and improves performance in low-intensity scenarios. Combining generative architectures with our fusion strategies is a promising direction, as generative models could reduce smoothing and peak underestimation at longer lead times while benefiting from the added atmospheric information provided by multi-modal datasets.
5.2 Variable Analysis
To quantify the individual contribution of each meteorological variable, we conduct a study in which each of the eight weather-station variables is removed while all other inputs remain unchanged. The ablation results in Table 4 provide further insight into how individual meteorological variables influence the performance of the two proposed models. Across both architectures, removing a single variable results in a decline in predictive skill, indicating that the station-derived features meaningfully complement the radar-based inputs. The magnitude of this decline reveals clear differences in the roles played by specific atmospheric variables.
For SmaAt-Krige-GNet which incorporates spatially Kriging interpolated weather shows the strongest sensitivity to humidity, atmospheric pressure, wind direction, and average wind speed. Removing any of these variables produces a noticeable degradation ($\Delta F1\approx 0.016–0.017$), corresponding to a contribution of roughly 2–3%. In contrast, variables that represent short-term fluctuations such as the standard deviations of wind speed and direction have less contribution to the model. The SmaAt-fUsion model shows a similar ranking of variable importance, though with generally smaller degradation when variables are removed. The reason is SmaAt-fUsion integrates station information only at the bottleneck rather than at multiple spatial scales. Nevertheless, the same four variables wind direction, atmospheric pressure, humidity, and average wind speed—remain the most influential. Removing these variables reduces F1 by approximately 0.011–0.015, corresponding to 1.4–1.9% contribution. The consistency of variable importance across both architectures indicates that our multi-source data fusion strategy effectively exploits the complementary strengths of weather-station variables and radar information.
@@ 6 Conclusion
This study introduced two novel deep learning models, SmaAt-fUsion and SmaAt-Krige-GNet, designed to improve precipitation nowcasting by integrating meteorological variables from weather stations alongside radar data. The experimental results demonstrate that both models have distinct advantages: SmaAt-Krige-GNet excels in low-intensity rainfall predictions by incorporating weather station data through Kriging-based spatial interpolation, while SmaAt-fUsion outperforms in high-intensity events by directly integrating station data into the network’s bottleneck. These findings highlight the potential of leveraging both discrete weather station data and spatially interpolated variables to improve deep learning-based precipitation forecasting. Future work includes extending the proposed fusion strategies to alternative architectures, additional multi-source meteorological datasets, and longer lead times to further assess their generality and robustness across diverse nowcasting settings.